# Structure-aware Domain Knowledge Injection for Large Language Models

### v2

@@ Abstract
This paper introduces a pioneering methodology, termed StructTuning, to efficiently transform foundation Large Language Models (LLMs) into domain specialists. It significantly reduces the training corpus requirement to a mere 0.3%, while achieving an impressive 50% of traditional knowledge injection performance. Our method is inspired by the educational processes of human students, particularly how structured domain knowledge from textbooks is assimilated and subsequently applied to tackle real-world challenges through specific exercises. Based on this, we propose a novel two-stage strategy for knowledge injection and alignment: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT). In the SCPT phase, we automatically extract the domain knowledge taxonomy and reorganize the training corpora, enabling LLMs to effectively link textual segments to targeted knowledge points within the taxonomy. In the SSFT phase, we explicitly prompt models to elucidate the underlying knowledge structure in their outputs, leveraging the structured domain insight to address practical problems. Our ultimate method has undergone extensive evaluations across model architectures and scales, using closed-book question-answering tasks on LongBench and MMedBench datasets. Remarkably, our method demonstrates the potential of comparable improvement against the state-of-the-art MMedLM2 on MMedBench, while significantly reducing the training costs to 5%. This breakthrough paves the way for scaling up our StructTuning for stronger domain-specific LLMs with comprehensive data utilization. Code is available at https://github.com/alibaba/struxgpt.
@@ 1 Introduction
Large language models (LLMs) have recently seen extensive deployment across various applications (Vaswani et al., 2017; Achiam et al., 2023; Jiang et al., 2023; Bi et al., 2024). When adapting foundational models (e.g., Llama series (Touvron et al., 2023a; b; Dubey et al., 2024)) for specialized AI assistants in distinct domains (Qiu et al., 2024; Guo et al., 2024), developers usually employ two techniques to enhance LLMs’ proficiency: retrieval-augmented generation (RAG) (Lewis et al., 2020) and domain knowledge injection (Gururangan et al., 2020). While RAG effectively utilizes an external knowledge base to augment information, the retrieval process’s inherent noise poses challenges to generating reliable responses, especially in scenarios requiring logical reasoning where there is a semantic gap between the user’s query and the knowledge base.(Zhang et al., 2023; Chen et al., 2023). Thus, another avenue of research focuses on injecting new knowledge into models via training techniques (Gu et al., 2021; Hu et al., 2021; Mecklenburg et al., 2024).
Continual pre-training (Sun et al., 2020; Ibrahim et al., 2024) has been preferred for integrating new, domain-specific knowledge into existing LLMs (Cui et al., 2023; Wang et al., 2023b; Qiu et al., 2024). Nevertheless, it often entails resource-intensive auto-regressive training on billions of tokens from the internet to learn fragmented knowledge points, rather than absorbing structured knowledge from a few domain-specific textbooks (Jin et al., 2020). For example, MMedLM (Qiu et al., 2024) curates 25.5B tokens to derive a medical model, and DeepSeek-Coder (Guo et al., 2024) uses 2T tokens for coding adaptation. The common failure to learn effectively from limited textbook content has been attributed to insufficient data diversity (Zhu & Li, 2023a), which however violates the observation during the human education process in Fig. 1: students gain knowledge by sequentially studying from textbooks, reviewing knowledge points and structures, and applying this knowledge through proper exercises. In this process, all the new data to learn are textbooks (structured content) and exercising examples (question-answering pairs), and students just adopt their world knowledge to memorize, understand, and apply the knowledge to become domain experts (Krathwohl, 2002; Yu et al., 2023).
@@
Figure 1: Discrepancy between human learning process and vanilla LLM adaptation paradigm. Human students learn structured knowledge through textbooks section by section, with particular exercises on related knowledge points. Traditional LLM adaptation involves continual pre-training on data chunks from randomly concatenated text segments, with aimless supervised fine-tuning for conversation alignment. The inherent property of structured knowledge is ignored.
@@
Inspired by this, we propose to inject the domain knowledge from textbooks into LLMs, as educating a human student, through a novel two-stage training strategy: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT).
@@
In the SCPT stage, we argue that high-quality textbook data can adequately infuse the domain knowledge (Gunasekar et al., 2023), where the organization of training corpora is crucial. In conventional paradigms, as illustrated in Fig. 1, text corpora are simply concatenated and divided into chunks of 2048 (Qiu et al., 2024) or 4096 (Guo et al., 2024), while the inherent structure of the texts (e.g., catalogs of textbooks) is disregarded. Instead, we propose an automatic approach to maintain each chunk’s knowledge structure. We view each chunk as a knowledge point, and employ advanced LLMs to efficiently extract domain knowledge taxonomy from the corpus, bypassing the need for manual annotation. Subsequently, LLMs are trained to predict the textual content (corresponding to the knowledge point) under the condition of the knowledge path within the domain structure, linking individual training chunks with the entire knowledge architecture. Finally, models are asked to memorize the knowledge structure to review the whole domain knowledge system.
@@
In the SSFT stage, the goal shifts from knowledge injection to enabling LLMs to recall and utilize their acquired knowledge to tackle real-world challenges. We explicitly elicit knowledge paths in LLMs’ responses, as a beacon for models to targeted information retrieval or logical reasoning for reliable responses. To this end, we derive a scalable strategy to generate question-answer pairs as practice exercises by powerful LLMs such as GPT4 (Achiam et al., 2023) or LLaMA3 (Dubey et al., 2024). In the scenarios with existing QA pairs like MMedBench (Qiu et al., 2024), we retrieve related knowledge structure and content, instructing LLaMA3 to provide explanations from questions to answers based on the knowledge paths. For datasets lacking specific QA samples like LongBench (Bai et al., 2023), we randomly select knowledge paths from the domain taxonomy and prompt LLaMA3 to craft question-answer-explanation triplets for training exercises.
@@
Our ultimate approach, termed StructTuning, outperforms conventional methods in domain knowledge injection by emulating human learning processes through SCPT and SSFT phases. We extensively evaluate StructTuning’s effectiveness across different model architectures and sizes. For domain-adapted language models, we first examine their capability to recall the injected knowledge through open-ended QA on the LongBench (Bai et al., 2023) dataset, then assess their application of this knowledge in addressing real-world issues via multiple-choice QA on MMedBench (Qiu et al., 2024). Both evaluations underscore the superiority of StructTuning. Remarkably, we achieve a 50% improvement in knowledge injection compared to the SOTA MMedLM2 in the medical domain, using merely 0.3% of the training data requirement. And StructTuning illustrates the potential of comparable performance with only 5% of training costs. These findings reveal the scalability of our method for enhancing domain-specific AI assistants with further comprehensive data utilization.
@@
  • We developed a scalable data construction framework to generate structure-aware training samples from original corpora, so as to facilitate the SCPT and SSFT stages.
@@
2 Related Work
Domain Adaptation for Large Language Models. While pre-trained LLMs possess promising capabilities, their performance is often hampered by the scope and recency of their training data, which particularly affects smaller models in downstream applications (Zhao et al., 2023; Wang et al., 2023a). Continual Pre-Training (CPT) addresses this by perpetually updating a pre-trained model with domain-specific content (Sun et al., 2020; Xu et al., 2023b). To keep pace with the latest information, models can be fine-tuned with supervised instruction-response pairs (SFT), thus staying current with the advancing knowledge landscape (Mecklenburg et al., 2024; Qiu et al., 2024). Existing literature confirms that combining CPT and SFT is effective for LLMs to remain precise and up-to-date in dynamic fields like law (Cui et al., 2023; Nguyen, 2023) and medicine (Wang et al., 2023b; Qiu et al., 2024). Our study builds upon this CPT-SFT framework, innovating with SCPT-SSFT strategies to efficiently and effectively infuse domain knowledge with the inherent structure hierarchy.
Structure-aware Knowledge Aggregation. A branch of researchers follows the conventional paradigm to extract entity-relation-entity triplets from texts to construct knowledge graphs (Pan et al., 2024), to enhance LLMs’s factual knowledge and logical reasoning by feature aggregation (Liu et al., 2020; Zhang et al., 2022), prompt engineering (Wen et al., 2023; Wang et al., 2023c), training data synthesis (Tang et al., 2024), etc. In these cases, each node corresponds to either a specific entity or an abstract concept, lacking the capability to present an informative and self-contained knowledge point. Some works have recently related a piece of descriptive text to a knowledge point, and constructed the knowledge structure for LLMs’ retrieval-augmented generation (Sarthi et al., 2024; Dong et al., 2024), where the top-to-down retrieval provides precise information-seeking paths along the knowledge structure. In this paper, we extend the structure-aware knowledge aggregation to LLMs’ training phase, injecting the whole domain knowledge structure into LLMs’ by linking the training samples into corresponding knowledge points and reasoning paths.
Data Augmentation and Synthesis. Traditional methods aim to artificially expand the training dataset size (Xu et al., 2023a; Mukherjee et al., 2023) or generate entirely new samples that could help models learn better or adapt to specific tasks (Tang et al., 2024). Yet, they often overlook the structured nature of domain knowledge, and the aimlessly generated samples may also lack diversity (Ovadia et al., 2023; Mecklenburg et al., 2024), leading to potentially suboptimal training outcomes when applied for domain adaptations (Mecklenburg et al., 2024; Tang et al., 2024). By contrast, our SSFT design is an innovative departure to address the challenge of retaining and utilizing the structured knowledge inherent in domain-specific content.
@@ 3 Methodology
Fig. 2 depicts our StructTuning methodology to inject domain knowledge into pre-trained LLMs using the inherent knowledge structure. With curated domain corpora (typically a few textbooks), we first develop an automatic approach to extract the knowledge structure, and associate text chunks to corresponding knowledge paths and points (Sec. 3.1). Then, we design a two-stage training strategy to inject the highly structured domain knowledge into language models by mimicking the human education process, comprising the SCPT (Sec. 3.2) and SSFT (Sec. 3.3) techniques.
Figure 2: Framework for structure-aware knowledge injection. We extract the inherent knowledge structure in the training corpus, and associate training chunks to corresponding knowledge points. Models are continually pre-trained on data chunks in the condition of the knowledge structure, and fine-tuned with supervised QA samples to elicit their learned knowledge to solve real-world questions (including knowledge-intensive (KI) QA, 2- or multi-hop QA, etc.).
3.1 Automatic Extraction of Knowledge Structure
@@
Then, we automatically aggregate knowledge points and extract the inherent structure hierarchy by leveraging advanced language models. Inspired by Liu et al. (2024a), we take the title list to instruct a specifically developed 7B model to identify the inherent knowledge structure within the text chunks. A prompt example is displayed in Fig. 3, and the detailed implementation is presented in Sec. B.1. In particular, Sec. B.1 and Sec. B.2 verify that our specialized 7B model can identify sufficiently precise knowledge structure for effective and efficient domain adaptation, as more powerful LLMs like LLaMA3-70B (Dubey et al., 2024) and GPT-3.5 (Brown et al., 2020) cannot bring significant enhancement while largely increase the inference costs.
Fig. 3 presents an example of the extracted structure, where we use the tree-like mindmap structure (Wen et al., 2023) to present the knowledge taxonomy from a textbook. And Fig. A1 showcases how to deal with non-textbook data. The whole process does not involve human annotation, which reduces the cost and makes our method scalable for larger domain training corpora.
Figure 3: Left: prompt template to extract hierarchical knowledge structures from the given content. Right: example of the extracted knowledge structure presented by a mindmap format.
After automatically extracting the domain knowledge structure and associating the original training chunks to related knowledge points, we delve into injecting domain knowledge through structure-aware continual pre-training (SCPT) and structure-aware supervised fine-tuning (SSFT).
@@ 3.2 Structure-aware Continual Pre-Training
In conventional knowledge injection methods, training corpora are randomly concatenated and chunked into text segments without distinguishing the original content, leading to the fact that models can only absorb domain knowledge that is emergent in the data diversity (Ovadia et al., 2023; Mecklenburg et al., 2024; Qiu et al., 2024). In this section, we present another solution to inject knowledge from limited pieces of textbooks by leveraging the highly abstractive and exhaustive domain knowledge structures for continual pre-training.
Figure 4: Example of prompt templates to bridge mindmap structure and textual contents.
We first transform the knowledge structure into natural languages using the same mindmap template in Fig. 3, and prepend it to each training chunk, forcing LLMs to memorize the textual content (knowledge points) in the condition of the associated knowledge path in the structure hierarchy. We collected 20 diversified templates from GPT-4 (Achiam et al., 2023) to bridge mindmap structures and training chunks, one of which is displayed in Fig. 4. The prepended mindmap, as well as the template, does not produce auto-regressive loss. Losses are only calculated in the content part. Formally, we turn the original language modeling in vanilla CPT to conditioned modeling (Keskar et al., 2019) in our SCPT stage:
$p({\bm{x}}^{k})=\prod_{i=1}^{n}{p(x^{k}_{i}|x^{k}_{<i})}\quad\Longrightarrow\quad p({\bm{x}}^{k}|{\bm{s}}^{k})=\prod_{i=1}^{n}{p(x^{k}_{i}|x^{k}_{<i},{\bm{s}}^{k})}$ (1)
where $p({\bm{x}}^{k})$ models the probability distribution for the $k$-th chunk ${\bm{x}}^{k}=(x^{k}_{1},\cdots,x^{k}_{n})$ via the chain rule of probability (Bengio et al., 2000) on each token $x^{k}_{i}$, and ${\bm{s}}^{k}$ denotes the associated knowledge mindmap paths.
As illustrated in Fig. 2 (b), after traversing the $m$ knowledge points in extracted structures, models are asked to recall the whole knowledge hierarchy, i.e., to model the composed probability distribution:
@@
Next, we will introduce how to teach LLMs to explicitly utilize their domain knowledge, which is learned in the SCPT stage, to solve practical problems by doing exercises with our SSFT technique.

### v3

@@ Abstract
This paper introduces a pioneering methodology, termed StructTuning, to efficiently transform foundation Large Language Models (LLMs) into domain specialists. It significantly reduces the training corpus needs to a mere 5% while achieving an impressive 100% of traditional knowledge injection performance. Motivated by structured human education, we propose a novel two-stage strategy for knowledge injection and alignment: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT). In the SCPT phase, we automatically extract the domain knowledge taxonomy and reorganize the training corpora, enabling LLMs to effectively link textual segments to targeted knowledge points within the taxonomy. In the SSFT phase, we explicitly prompt models to elucidate the underlying knowledge structure in their outputs, leveraging the structured domain insight to address practical problems. Our ultimate method was extensively evaluated across model architectures and scales on LongBench and MMedBench datasets, demonstrating superior performance against other knowledge injection methods. We also explored our method’s scalability across different training corpus sizes, laying the foundation to enhance domain-specific LLMs with better data utilization. Code is available at: https://github.com/alibaba/struxgpt.
Structure-aware Domain Knowledge Injection for Large Language Models
Kai Liu1,2111Equal contribution. Work done during Kai Liu’s research internship at Alibaba Cloud. Email: kail@zju.edu.cn.,  Ze Chen2111Equal contribution. Work done during Kai Liu’s research internship at Alibaba Cloud. Email: kail@zju.edu.cn.,  Zhihang Fu2222Corresponding authors. Email: zhihang.fzh@alibaba-inc.com, 0006236@zju.edu.cn.,  Wei Zhang1,  Rongxin Jiang1, Fan Zhou1222Corresponding authors. Email: zhihang.fzh@alibaba-inc.com, 0006236@zju.edu.cn.,   Yaowu Chen1,  Yue Wu2,  Jieping Ye2 1Zhejiang University,   2Alibaba Cloud
@@ 1 Introduction
Figure 1: Discrepancy between human education and vanilla LLM adaptation. Human students learn structured knowledge through textbooks section by section, with particular exercises on related knowledge points. Traditional LLM adaptation continually pre-trains on data chunks from randomly concatenated text segments, with aimless supervised fine-tuning for conversation alignment. The inherent knowledge structure is ignored.
Large language models (LLMs) have recently seen extensive deployment across various applications (Vaswani et al., 2017; Achiam et al., 2023; Jiang et al., 2023; Bi et al., 2024). When adapting foundational models (e.g., Llama series (Touvron et al., 2023a, b; Dubey et al., 2024)) to specialized AI assistants in distinct domains, developers usually employ two techniques to enhance LLMs’ proficiency: retrieval-augmented generation (RAG) (Lewis et al., 2020) and domain knowledge injection (Gururangan et al., 2020). While RAG effectively utilizes an external knowledge base to augment information, the retrieval process’s inherent noise poses challenges to generating reliable responses, especially in scenarios requiring logical reasoning where there is a semantic gap between the user’s query and the knowledge base.(Zhang et al., 2023; Chen et al., 2023). Thus, another avenue tries to inject new knowledge to LLMs via training techniques (Gu et al., 2021; Hu et al., 2021; Mecklenburg et al., 2024).
Continual pre-training (Sun et al., 2020; Ibrahim et al., 2024) is widely used for injecting domain-specific knowledge (Cui et al., 2023; Wang et al., 2023b; Qiu et al., 2024). However, it often requires training on billions of internet tokens to capture fragmented knowledge, rather than leveraging a few structured textbooks (Jin et al., 2020). For instance, MMedLM (Qiu et al., 2024) uses 25.5B tokens for medical modeling, while DeepSeek-Coder (Guo et al., 2024) processes 2T tokens for coding adaptation. The limited ability to learn effectively from textbooks was attributed to insufficient data diversity (Zhu and Li, 2023a), which however violates the observation during the human education process in Fig. 1: students gain knowledge by sequentially studying from textbooks, reviewing knowledge points and structures, and applying this knowledge through proper exercises. Here, all the new data to learn are textbooks (structured content) and exercising examples (question-answering pairs), and students just adopt their world knowledge to memorize, understand, and apply the knowledge to become domain experts (Krathwohl, 2002; Yu et al., 2023).
@@
As educating human students, we propose to inject structured domain knowledge into LLMs via two steps: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT).
@@
In the SCPT stage, we argue that high-quality textbook data (as well as regular corpora from the Internet) can adequately infuse domain knowledge (Gunasekar et al., 2023), where the organization of training corpora is crucial. In conventional paradigms (Fig. 1), text corpora are simply concatenated and divided into chunks of 2048 (Qiu et al., 2024) or 4096 (Guo et al., 2024) tokens, while the inherent semantic structure (e.g., catalogs of textbooks) is discarded. Instead, we view each chunk as a knowledge point and automatically extract domain knowledge taxonomy from the whole corpus. Subsequently, LLMs are trained to predict textual content (corresponding to a knowledge point) under the condition of the knowledge path within the domain structure, linking individual training chunks with the entire knowledge architecture. Finally, models have to memorize the entire structure to review the whole domain knowledge system.
@@
In the SSFT stage, the goal shifts from knowledge injection to enabling LLMs to recall and utilize their acquired knowledge to tackle real-world challenges. We explicitly elicit knowledge paths in LLMs’ responses, as a beacon for models to targeted information retrieval or logical reasoning for reliable responses. To this end, we derive a scalable strategy to generate question-answer pairs as practice exercises by open-sourced LLMs or API, such as LLaMA3 (Dubey et al., 2024) and GPT4 (Achiam et al., 2023). In the scenarios with existing QA pairs like MMedBench (Qiu et al., 2024), we retrieve the related knowledge structure and content, instructing LLaMA3 to provide explanations from questions to answers based on the knowledge paths. For datasets lacking specific QA samples like LongBench (Bai et al., 2023b), we randomly select knowledge paths from the domain taxonomy and prompt LLaMA3 to craft question-answer-explanation triplets for training exercises.
@@
Our ultimate approach StructTuning has been extensively evaluated across different model architectures and sizes. In particular, we first examine their capability to recall the knowledge injected through open-ended QA on the LongBench (Bai et al., 2023b) dataset, then assess the application of injected knowledge to address real-world issues through multiple-choice QA on MMedBench (Qiu et al., 2024). Both evaluations underscore the superiority of StructTuning, surpassing other SOTA domain knowledge injection methods (Cheng et al., 2023; Zhang et al., 2024). Remarkably, we achieve a 50% improvement in knowledge injection compared to SOTA MMedLM2 in the medical domain, using only 0.3% of the training data requirement. Furthermore, StructTuning exhibits good scalability, achieving comparable performance with only 5% of the training data. These findings reveal our superiority in enhancing domain-specific AI assistants with more efficient data utilization.
Figure 2: Framework for structure-aware knowledge injection. We extract the inherent knowledge structure in the training corpus, and associate training chunks to corresponding knowledge points. Models are continually pre-trained on data chunks in the condition of the knowledge structure, and fine-tuned with supervised QA samples to elicit their learned knowledge to solve knowledge-intensive (KI) and 2- or multi-hop questions in the real world.
@@
  • We developed a scalable data construction framework to generate structure-aware training samples from original corpora to facilitate the SCPT and SSFT stages.
@@
2 Related Works
Here we briefly discuss the closely related works. A detailed discussion can be found in Appendix C.
Domain Adaptation for LLMs. To address the domain adaptation problem, a pre-trained model will be continually pre-trained (CPT) with domain-specific content (Sun et al., 2020; Xu et al., 2023b), and fine-tuned with supervised instruction-response pairs (SFT) to keep advancing interactive capabilities (Mecklenburg et al., 2024; Qiu et al., 2024). This paradigm is validated efective in dynamic fields like medicine (Wang et al., 2023b; Qiu et al., 2024) and coding (Roziere et al., 2023; Guo et al., 2024). Our study builds upon this CPT-SFT framework, innovating with SCPT-SSFT strategies to efficiently and effectively infuse domain knowledge with the inherent structure hierarchy.
Structure-aware Knowledge Aggregation. In conventional paradigms, researchers extract entity-relation-entity triplets from texts to construct knowledge graphs (Pan et al., 2024), to enhance LLMs’s factual knowledge and logical reasoning(Zhang et al., 2022; Wen et al., 2023). Here, each node corresponds to either a specific entity or an abstract concept, lacking the capability to present an informative and self-contained knowledge point. This paper extends to structure-aware knowledge aggregation on existing training corpora, injecting the whole domain knowledge structure into LLMs’ by linking training samples to corresponding knowledge points and reasoning paths.
Data Augmentation and Synthesis. Traditional methods aim to artificially expand the training dataset size (Xu et al., 2023a; Mukherjee et al., 2023) or generate entirely new samples to adapt LLMs to specific tasks (Tang et al., 2024). Yet, they often overlook the structured nature of domain knowledge, while the aimlessly generated samples may also lack diversity (Ovadia et al., 2023; Mecklenburg et al., 2024) and cannot cover the domain knowledge points (Mecklenburg et al., 2024; Tang et al., 2024). By contrast, our SSFT design is an innovative departure to address the challenge of retaining and utilizing the structured knowledge inherent in domain-specific content.
@@ 3 Methodology
Figure 3: Left: extracted knowledge structure. Right: template to bridge mindmap structure and textual contents.
Fig. 2 depicts our StructTuning methodology to inject domain knowledge into pre-trained LLMs using the inherent knowledge structure. With curated domain corpora (typically a few textbooks), we first extract the knowledge structure, and associate text chunks to corresponding knowledge paths and points (Sec. 3.1). Then, we design a two-stage training strategy to inject the highly structured domain knowledge into language models by mimicking the human education process, comprising the SCPT (Sec. 3.2) and SSFT (Sec. 3.3) techniques.
3.1 Knowledge Structure Extraction
@@
Then, we aggregate knowledge points and extract the inherent structure hierarchy by leveraging advanced language models. Inspired by Liu et al. (2024a), we take the title list to instruct a specifically developed 7B model to identify the inherent knowledge structure (as exemplified in Fig. 3) within the text chunks, and Fig. A1 showcases how to deal with non-textbook data. The whole process does not involve human annotation, which reduces the cost and makes our method scalable for larger domain training corpora.
In particular, Sec. B.1 and Sec. B.4 verify that our specialized 7B model can identify sufficiently precise knowledge structure for effective and efficient domain adaptation, as more powerful LLMs like LLaMA3-70B (Dubey et al., 2024) and GPT-3.5 (Brown et al., 2020) cannot bring significant enhancement while largely increase the inference costs.
@@ 3.2 Structure-aware Continual Pre-Training
In conventional knowledge injection, training corpora are randomly concatenated and chunked into text segments without distinguishing the original content, making models only able to absorb domain knowledge emerging in data diversity (Ovadia et al., 2023; Mecklenburg et al., 2024; Qiu et al., 2024). In this section, we present another solution to inject knowledge from limited text corpora by leveraging the highly abstractive and exhaustive domain knowledge structures for continual pre-training.
We first transform the knowledge structure into natural languages using the same mindmap template Wen et al. (2023) in Fig. 3 (left), and prepend it to each training chunk, forcing LLMs to memorize the textual content (knowledge points) in the condition of the associated knowledge path in the structure hierarchy. 20 diversified templates (see Fig. A5) are collected from GPT-4 (Achiam et al., 2023) to bridge mindmap structures and training chunks, one of which is displayed in Fig. 3 (right). The prepended mindmap, as well as the template, does not produce auto-regressive loss. Losses are only calculated in the content part. Formally, we turn the original language modeling in vanilla CPT to conditioned modeling (Keskar et al., 2019) in our SCPT stage:
$p({\bm{x}}^{k})=\prod_{i=1}^{n}{p(x^{k}_{i}|x^{k}_{<i})}\quad\Longrightarrow$ (1)
$p({\bm{x}}^{k}|{\bm{s}}^{k})=\prod_{i=1}^{n}{p(x^{k}_{i}|x^{k}_{<i},{\bm{s}}^{k})}$
where $p({\bm{x}}^{k})$ models the probability distribution for the $k$-th chunk ${\bm{x}}^{k}=(x^{k}_{1},\cdots,x^{k}_{n})$ via the chain rule of probability (Bengio et al., 2000) on each token $x^{k}_{i}$, and ${\bm{s}}^{k}$ denotes the associated knowledge mindmap. Sec. B.5 extensively investigates the effectiveness of our SCPT strategy.
As Fig. 2 (b) shows, after traversing the $m$ knowledge points in extracted structures, models are asked to recall the whole knowledge hierarchy, i.e., to model the composed probability distribution: